### v1

# Weichuang Shao Iman Yi Liao Tomas Henrique Bode Maul Tissa Chandesa

Abstract
Recent foundational models, SSAST, EAT, HuBERT, Qwen-Audio, and Audio Flamingo, achieve top-tier results across standard audio benchmarks but are limited by fixed input rates and durations, hindering their reusability. This paper introduces the Augmentation-driven Multiview Audio Transformer (AMAuT), a training-from-scratch framework that eliminates the dependency on pre-trained weights while supporting arbitrary sample rates and audio lengths. AMAuT integrates four key components: (1) augmentation-driven multiview learning for robustness, (2) a conv1 + conv7 + conv1 one-dimensional CNN bottleneck for stable temporal encoding, (3) dual CLS + TAL tokens for bidirectional context representation, and (4) test-time adaptation/augmentation ($TTA^{2}$) to improve inference reliability. Experiments on five public benchmarks, AudioMNIST, SpeechCommands V1 & V2, VocalSound, and CochlScene, show that AMAuT achieves accuracies up to 99.8% while consuming less than 3% of the GPU hours required by comparable pre-trained models. Thus, AMAuT presents a highly efficient and flexible alternative to large pre-trained models, making state-of-the-art audio classification accessible in computationally constrained settings.
Keywords: Transformer, Audio Classification, Multiview Augmentation, Test-time Adaptation, Test-time Augmentation
1 Introduction
(a) Training and TTDA topology. AMAuT supports more than two views using similar processing steps. Each view processing is independent before combining the loss.
(b) Prediction topology.
Figure 1: Overall architecture of the proposed AMAuT framework. Here, C denotes the number of classes, while N, F, T, and K depend on the dataset.
Audio classification is a machine learning task that aims to identify semantic categories from audio signals. Pre-trained models such as SSAST [16], EAT [6], HuBERT [19], Qwen-Audio [8], and Audio Flamingo [25], have achieved state-of-the-art performance on benchmark datasets including SpeechCommands V1 (SC1) [42], SpeechCommands V2 (SC2) [42], AudioMNIST (AM) [4], VocalSound (VS) [17, 15], and CochlScene (CS) [21]. However, these models rely heavily on pre-training to reach such performance levels, which imposes several practical limitations. Specifically, reusing pre-trained weights from these models requires a fixed sample rate (16 kHz for SSAST, EAT, HuBERT, and Qwen-audio, and 44.1 kHz for Audio Flamingo) and specific audio length (1 s, 5 s, or 10 s). These constraints restrict model reusability and make it difficult to apply the models to new datasets with different sample rates or audio lengths.
Although resampling techniques can be applied to meet the required rates, such preprocessing often introduces information loss when downsampling or artificial data when upsampling. For instance, [29] reported that the main frequency of bird sounds is below 12 kHz, implying that, based on the information theory [39] (Nyquist theorem), the sample rate should exceed 24 kHz to preserve information. However, 24 kHz is not a standard rate supported by most existing pre-trained models. Furthermore, as audio recording hardware continues to evolve, producing higher-quality data at higher sample rates, a fixed sample rate methodology cannot take advantage of such an improvement in the recording device.
The same limitation applies to audio length. SSAST and EAT, for example, support only predefined durations of 1, 5, or 10 seconds [16, 6], while Audio Flamingo allows a maximum of 33.25 seconds [25]. When the input audio length or sample rate does not match the model’s configuration, pre-training must be repeated or the audio must be forcibly adjusted, both of which are computationally expensive.
Consequently, this paper addresses a complementary challenge: how to eliminate the fixed input constraints of pre-trained audio transformers. To answer this question, this study proposes the Augmentation-driven Multiview Audio Transformer (AMAuT), a training-from-scratch framework (see Figure 1) that eliminates the dependency on pre-trained weights. This design offers two major advantages: (1) AMAuT accepts arbitrary input sample rates and audio lengths, thereby increasing flexibility and adaptability. (2) AMAuT significantly reduces computational requirements (GPU hours) by avoiding large-scale pre-training. To maintain performance, AMAuT leverages test-time adaptation/augmentation ($TTA^{2}$) to enhance inference robustness.
Major contributions of this paper include:
  1. 1.
    Augmentation-driven multiview learning: a training strategy that simultaneously optimizes multiple augmented versions of each audio sample to enhance generalization.
  2. 2.
    Flexible 1D CNN bottleneck: a conv1 + conv7 + conv1 design that converts variable-length audio inputs into uniform embeddings without resampling.
  3. 3.
    Dual CLS + TAL tokens: a novel token pairing that captures bidirectional contextual information within the Transformer encoder.
  4. 4.
    Test-time adaptation/augmentation ($TTA^{2}$): an inference strategy combining entropy-based domain adaptation and ensemble-style augmentation for improved robustness.
Comprehensive evaluations across five benchmarks reveal that AMAuT achieves accuracy comparable to or surpassing pre-trained counterparts while requiring less than 3% of their GPU hours. This efficiency highlights the potential of pretrain-free, adaptable Transformers for real-world audio classification where computational budgets are limited.
2 Related Works
2.1 Transformer-based Audio Models
The transformer architecture, initially developed for language processing, has been effectively adapted for vision and audio tasks. Vision Transformer (ViT) [11] and Distilled ViT (DeiT) [35] operate by dividing images into fixed-size patches (e.g., $16\times 16$ or $32\times 32$), a concept later transferred to the audio domain, such as Audio Spectrogram Transformer (AST) [14]. AST represents audio as Mel-spectrograms [36, 20], turning temporal signals into two-dimensional time–frequency grids before applying patch-based Transformer processing with fixed-size patches ($16\times 16$). While AST reuses ImageNet-trained weights [9, 14], SSAST [16] processes pre-training on AudioSet [13] and LibriSpeech [30] rather than reusing ImageNet-trained weights for extending AST.
Models such as Wav2Vec 2.0 [3] and HuBERT [10] process raw wavforms using 1D CNN-Transformer architectures. Wav2Vec 2.0 is composed of seven convolutional layers equipped with LayerNorm and GELU activations, and its large-scale pre-training on LibriSpeech [30] and Libri-Light [22] reportedly required about 7.5 days on 128 GPUs. HuBERT follows a similar CNN-Transformer architecture to Wav2Vec 2.0 but extends it with a BERT-style [10] masked prediction mechanism that integrates a waveform encoder. Like Wav2Vec 2.0, AMAuT adopts a 1D CNN-Transformer structure but differs fundamentally: AMAuT trains from scratch without pre-training, uses 1D CNNs on Mel-spectrogram horizontal tokens, and replaces LayerNorm + GELU with BatchNorm + ReLU. These changes improve both efficiency and accuracy (99.82% vs. 99.4% on AudioMNIST) between Wav2Vec 2.0 and AMAuT.
The spectrogram transformers [43] contributed a Tempo-Frequency sequential (TFS) attention for separating mel-spectrogram outputs within frequency direction instead of $16\times 16$ regions in AST [14] that are similar to the horizontal tokenization of AMAuT (see subsection 3.3). However, [43] inherits the methodology of AST and ViT, and then [43] employs a different model architecture, as it incorporates 2D CNN layers.
Overall, both AST and SSAST constrain input to fixed sample rates (typically 16 kHz) and durations (1 s, 5 s, or 10 s) when reusing their pre-trained weights; Wav2Vec 2.0 and HuBERT are computationally expensive and also fixed at 16 kHz. AMAuT removes these constraints, allowing flexible sample rates and audio lengths while matching or exceeding the accuracy of pre-trained models at a cost that is at most 3% of their GPU hours.
2.2 Test-Time Domain Adaptation
Test-time domain adaptation (TTDA) only uses the test set for performance improvement by adapting (partial) model parameters of a trained model from a source domain to a target domain [28]. In AMAuT, this is achieved through a composite loss function $\mathcal{L}_{TTDA}$ that combines three complementary objectives: Nuclear-Norm Maximization, Entropy Minimization, and Generalized Entropy (see Eq.7, Eq.8, and Eq.9 in Appendix A). These components have been applied separately in previous studies, such as CoNMix [26] for test-time domain adaptation, TENT [41] for entropy-based test-time batch adaptation, and SGEM [23] for Automatic Speech Recognition, but AMAuT unifies them for the first time. This integration allows efficient test-time domain adaptation and improves robustness under moderate distribution shifts without additional supervision.
3 Methodology
3.1 Overview
The proposed Augmentation-driven Multiview Audio Transformer (AMAuT) architecture (see Figure 1) follows a hybrid CNN-Transformer pipeline designed for flexibility and efficiency. Unlike prior models constrained by pre-trained feature extractors, AMAuT is trained entirely from scratch and can ingest audio at arbitrary sampling rates and lengths.
During training, each audio sample undergoes multiview augmentation, producing several perturbed variants (time-shifted, noise-added, or background-mixed). These parallel views are passed through identical processing streams, and their corresponding losses are jointly optimized to encourage invariance to perturbations.
Each augmented view is first converted to a Mel-spectrogram, then segmented along the temporal dimension through horizontal tokenization. A 1D CNN transforms these tokens into a fixed-size representation (768 × K), which captures fine-grained temporal dependencies with reduced computational cost compared to 2D CNNs.
The resulting embeddings enter the vertical embedding transformer, including CLS + TAL tokens and learnable positional embeddings. Multi-head self-attention layers refine the token relationships before classification.
During adaptation, AMAuT applies $TTA^{2}$, a lightweight refinement stage combining (a) test-time domain adaptation (TTDA) that updates internal statistics using unlabeled test data, and (b) test-time augmentation (TTAu) by averaging predictions across multiple augmented inputs. Together, these mechanisms improve robustness to unseen conditions without additional training labels.
3.2 Multiview Augmentation
During both training and test-time domain adaptation (TTDA) (see Figure 1(a)), the AMAuT framework employs augmentation-driven multiview learning to improve model robustness.
Multiple augmented views are generated from each input sample, and each view contributes to the overall loss calculation (see Algorithm 1 in Appendix A). Rather than computing gradients for each loss individually, AMAuT aggregates all losses before performing gradient updates:
$\sum^{A}_{i}\mathcal{L}(f(x_{i};\theta))$ (1)
where $\mathcal{L}$ denotes the loss function, $f(\cdot,\theta)$ represents the AMAuT model, $A$ denotes the number of augmentation-driven views, and $\{x_{i}|1\leq i\leq A\}$ are the augmented views of the same audio sample $x$.
During validation, AMAuT uses the original audio input $x$ without multiview augmentation to ensure a fair comparison with other state-of-the-art models (see Figure 1(b)).
In the experiments, four augmented views were used during training and two during TTDA (see Subsection 3.7). All augmentations introduce random variations, such as time shifts (Figure 1(a) demonstrates the left and right time-shift augmentations) and Gaussian or background noise, ensuring that each epoch encounters unique augmented samples.
Specifically:
  • Training (four augmentations): (1) random left or right time shift (0-17%), (2) Gaussian noise (noise-to-audio ratio = 0.015, and regenerate noise each time), and (3-4) random clips from two background noises (“dude miaowing” and “pick noise” from SpeechCommands [42]) with an SNR of 50 dB.
  • TTDA (two augmentations): (1) random right time shift (0-17%), and (2) random left time shift (0-17%).
3.3 Spectrogram and Horizontal Tokenization
Before tokenization, each audio sample is transformed into a 2D Mel-spectrogram, where the height and width correspond to frequency and time dimensions, respectively. AMAuT processes only mono audio. For multichannel audio, it converts the signals to mono by averaging across channels as follows:
[…]
where $x=\{x_{i,j}|1\leq i\leq\mathcal{C},1\leq j\leq N\}$ is a multi-channel audio signal, and $\text{sterToMono}(x)$ is the resulting mono signal. Here, $\mathcal{C}$ represents the number of channels and $N$ is the length of each channel.
After conversion to mono, the spectrogram has a shape of $1\times F\times T$, where $F$ and $T$ denote the frequency and time dimensions, respectively. The channel dimension (1) is removed, reshaping the spectrogram to $F\times T$ for subsequent tokenization (see Figure 1).
Horizontal tokenization divides the spectrogram along the time dimension, producing a sequence of temporal tokens. These tokens serve as input to the subsequent 1D CNN module, which efficiently captures temporal dependencies.
3.4 1D CNNs
The architecture uses 1D CNNs because it interprets the spectrogram’s frequency bins as feature channels, leveraging the independent nature of different frequencies in an audio signal. Further, 1D CNNs consume less GPU memory and computation cost, as a 1D convolution (conv7) has a smaller kernel size than a 2D convolution (conv3×3). The latter requires two-dimensional kernels that significantly increase GPU memory usage. The 1D CNN module processes the horizontal tokens and reshapes them into a consistent feature size of $768\times K$ (see Figure 1).
Table 1: List the AMAuT configurations for different datasets, including the number of convolutional layers, GPU memory usage, and total parameters.
Set CNN Depth GPU Memory Parameters
AM 49 2.63 GB 99.1 M
SC1 61 4.79 GB 103.1 M
SC2 61 4.20 GB 103.1 M
VS 64 9.47 GB 97.0 M
CS 79 12.89 GB 101.0 M
Longer audio inputs require deeper CNN stacks to reduce temporal resolution to 768 tokens, resulting in dataset-specific CNN configurations (see Table 1).
Figure 2: Structure of the 1D CNN module used in AMAuT
As illustrated in Figure 2, the 1D CNN module consists of:
  • Tail block: a conv14 + max-pooling layer for dimensionality reduction;
  • Bottleneck blocks (yellow): multiple conv1 + conv7 + conv1 sequence blocks for feature extraction;
  • Head block: a final conv1 layer that fixes the output shape to be $768\times K$.
This architecture efficiently extracts temporal representations while maintaining manageable computational complexity, making it suitable for diverse audio lengths and sample rates.
3.5 Vertical Embedding
Figure 3: Architecture of the Vertical Embedding module in AMAuT
As shown in Figure 3, the vertical embedding module receives 1D CNN features with a shape of $768\times K$. Since vertical embedding converts the tensor to $K\times 768$ by transpose, the process is referred to as vertical embedding.
CLS+TAL Tokens and Learnable Positional Embedding
AMAuT introduces two learnable tokens: the class token (CLS) and the tail token (TAL) (see Eq. 3 and Figure 3). Unlike conventional designs such as DeiT [35], where CLS and distillation tokens are placed adjacently, AMAuT positions CLS and TAL at opposite ends of the token sequence. This design assumes that tokens at different spatial positions may encode distinct contextual information during transformer processing.
Before the attention extraction stage, AMAuT applies a learnable positional embedding and a random dropout layer (15% drop rate) to the concatenated token sequence:
$h$ $=\text{concate}(\text{CLS},x,\text{TAL})$ (3)
$\hat{x}$ $=\text{mask}(h+\mathcal{P})$ (4)
where $x$ denotes the tokenized features, $\hat{x}$ the embedding input, $\mathcal{P}=\{\mathcal{P}_{0},\cdots,\mathcal{P}_{768+1}\}$ the positional embedding table, and $\text{mask}(\cdot)$ the dropout layer.
Attention Extraction
This component follows the Transformer [38] architecture and comprises multiple transformer blocks. Two important configurations are adopted:
  1. 1.
    The dropout rate for both multi-head attention and MLP is set to 0%, consistent with models such as SSAST [16], AST [14], and CoNMix [26].
  2. 2.
    AMAuT uses a full attention matrix rather than a lower-triangular one, as it performs classification rather than generative prediction. This choice enhances performance by preserving all temporal dependencies.
3.6 Classifier
Figure 4: Classifier architectures used in AMAuT for different audio durations. (a) For audio >>1s, only CLS + TAL tokens are used, followed by a 1D CNN and a fully connected (FC) layer. (b) For 1s inputs, using all features with AvgPool and FC layers.
As illustrated in Figure 4, AMAuT employs two classifier architectures based on the input audio duration.
  • For inputs longer than 1 second (Figure 4(a)):
    Only the CLS and TAL tokens are used. The classifier comprises one 1D CNN layer followed by a fully connected (FC) layer that maps features to class probabilities.
  • For 1-second inputs (Figure 4(b)):
    All transformer-extracted features are used. The classifier applies average pooling, followed by three FC layers with batch normalization, reducing the dimensionality to the number of output classes.
This dual-strategy design ensures efficient processing for short inputs while maintaining representational richness for longer sequences.
3.7 Test-time Adaptation / Augmentation ($TTA^{2}$)
At inference, AMAuT employs $TTA^{2}$, combining two complementary strategies:
Test-time Doman Adaptation (TTDA)
This subsection investigates an uncommon application of Test-time Domain Adaptation (TTDA): improving accuracy when training and test distributions are similar. Typically, TTDA is employed to address distribution shifts, but we explore its utility even when such a shift is minimal [28].
TTDA enhances performance by adapting a trained model to the test data distribution without requiring labels [28]. During TTDA, all test features are visible, but labels remain hidden. The model updates its internal parameters to better align with the target domain before final prediction.
TTDA follows the same computational structure as training but employs a different loss $\mathcal{L}_{TTDA}$ rather than the training loss $\mathcal{L}_{Tr}$ (see Figure 1(a)). AMAuT combines three entropy-based losses into a single objective (see Eq.7, Eq.8, and Eq.9 in Appendix A):
$\mathcal{L}_{TTDA}=\alpha\mathcal{L}_{NM}+\beta\mathcal{L}_{EN}+\gamma\mathcal{L}_{GEN}$ (5)
where $\alpha$, $\beta$, and $\gamma$ are scaling coefficients (see Table11 in Appedix B).
By integrating these complementary objectives, AMAuT improves inference robustness in scenarios where training and testing data share similar but not identical distributions.
Test-time Augmentation (TTAu)
To further enhance inference robustness, AMAuT adopts test-time augmentation (TTAu) [24]. Unlike TTDA, TTAu does not update model parameters; instead, it refines predictions by averaging outputs from multiple augmented versions of the same input. Three TTAu strategies are evaluated: Augmentation-driven refinement (Aug), Multi-training-based refinement (Mlt), and Hybrid refinement (Hyb) (see Eq.10, Eq.11, and Eq.12 in Appendix A, separately).
3.8 Learning Rate and Optimization
Both the training loss $\mathcal{L}_{Tr}$ (see Eq. 13 in Appendix A) and TTDA loss $\mathcal{L}_{TTDA}$ (see Eq. 5) employ a dynamic learning rate scheduler to accelerate convergence. The scheduler begins with a relatively large learning rate and gradually reduces it as training progresses, ensuring stable optimization. The update rule is defined as
$lr_{e}=\begin{cases}lr_{e-1}(1+\lambda\cdot e/\eta)^{0.75},&\text{if }e<\eta\\ lr_{e-1},&\text{otherwise}\end{cases}$ (6)
where $e$ denotes the epoch number, $10\leq\lambda\leq 40$ controls the rate of change, $200\geq\eta\geq 40$ defines the cardinality. See more details of hyperparameter settings from Table 12 in Appendix B.
During both training and TTDA, AMAuT uses the stochastic gradient descent (SGD) optimizer with weight decay $10^{-3}$, momentum 0.9, and Nesterov acceleration enabled.
4 Experiments and Results
AMAuT integrates several architectural and training strategies, including augmentation-driven multiview learning, conv1 + conv7 + conv1 CNN bottlenecks, flexible handling of arbitrary sample rates and audio lengths, CLS+TAL learnable positional embeddings, and $TTA^{2}$. Collectively, these components enable AMAuT to achieve strong performance when trained entirely from scratch, without relying on large-scale pre-training.
4.1 Datasets
Table 2: Dataset Properties
Set Size Length Rate
AM 30000 1 s 48 kHz
SC1 64721 1 s 16 kHz
SC2 105829 1 s 16 kHz
VS 20977 12 s 16 kHz
CS 76115 10 s 44.1 kHz
To evaluate the proposed AMAuT framework, this study employed five publicly available benchmark datasets: AudioMNIST (AM) [4], SpeechCommands V1 (SC1) [42], SpeechCommands V2 (SC2) [42], VocalSound (VS) [17, 15], and CochlScene (CS) [21].
As shown in Table 2, these datasets cover a diverse range of sampling rates and audio durations, enabling a comprehensive analysis of AMAuT’s flexibility and robustness. Specifically, AM operates at 48 kHz, CS at 44.1 kHz, and the remaining datasets at 16 kHz. Additionally, audio lengths also vary substantially, including 12 s for VS, 10 s for CS, and 1s for the others, allowing AMAuT to be evaluated on both short and long clips. Thus, this experimental configuration enabled consistent comparisons across all datasets.
For model training, AMAuT used the training and validation splits for AM, and the training sets for SC1, SC2, VS, and CS.
4.2 Ablation Studies
Effectiveness of Multiview Augmentation
The experiments support the claim that augmentation-driven multiview learning provides superior performance to conventional augmentation methods. Unlike conventional sequential or single-view augmentation, the proposed multiview approach applies multiple random augmentations to each input sample during both training and test-time domain adaptation (TTDA). Rather than generating and storing augmented data offline, AMAuT performs augmentations in real time, enabling transformations to vary across epochs. Consequently, the model is continually exposed to diverse variants of the same source sample, improving generalization and robustness.
Table 3: Comparison of test accuracies between sequential augmentation and augmentation-driven multiview learning on CochlScene (CS) and SpeechCommands V1 (SC1) before applying $TTA^{2}$. “1-Aug” and “4-Aug” indicate the use of one or four augmented views, respectively. The arrows ($\uparrow$, $\downarrow$) denote performance improvement or degradation relative to the baseline.
Set Aug-method 1-Aug(%) 4-Aug(%)
[…]
Compared with the traditional dataset-extension method (offline), where augmented samples are pre-generated and stored for later training, the augmentation-driven multiview learning approach offers two significant advantages:
  1. 1.
    Dynamic sample variation across epochs:
    Because augmentation is performed on-the-fly and randomly for each sample, the augmented versions differ each epoch. This continual variation prevents the model from overfitting to a fixed augmented dataset and enhances its exposure to a wider range of feature perturbations.
  2. 2.
    Joint optimization of related augmented samples:
    All augmented views originating from the same input are simultaneously passed through the network during training. This enables AMAuT to recognize that these views share a common semantic source but contain different biases, allowing the model to learn invariances more effectively. As a result, the model gains improved robustness against real-world distortions such as time shifts, Gaussian noise, and background interference.
Together, these two advantages enhance AMAuT’s generalization ability over offline augmentation methods.
For online augmentation methods, as shown in Table 3, the proposed multiview approach achieved higher test accuracies on both datasets, for example, an increase from 65.37% to 68.87% on CS and from 95.26% to 95.93% on SC1, demonstrating that real-time, multiview training substantially improves performance without additional pre-training or data storage requirements.
CNN Kernel Design
In many existing CNN-based audio classification architectures, relatively small convolutional kernels are commonly used. For example, HuBERT [19] and Wav2Vec 2.0 [3] employ bottleneck structures such as conv2 + conv3, while other hybrid CNN-Transformer architectures like ENACT-Heart [18] and CNN-TE [5] use kernel sizes of 3, 2, and 1. These designs typically include only one or two large convolutional layers, such as conv14 or conv16, at the head or tail of the CNN stack, rather than throughout its body.
In contrast, this study demonstrates that adopting a larger kernel size, specifically conv7, provides a better trade-off between efficiency and performance for both 1-second and 10-second audio datasets. The conv7 kernel allows AMAuT to capture longer temporal dependencies without significantly increasing computational complexity.
Table 4: Performance comparison between conv3 and conv7 kernel structures on CochlScene (CS) and SpeechCommands V1 (SC1) before applying $(TTA)^{2}$, evaluated on a single RTX 4090 GPU. “Length” indicates audio length in seconds, and “1-Aug” / “4-Aug” denote the number of augmented views used during testing.
Set Length Structure 1-Aug(%) 4-Aug(%)
[…]
As presented in Table 4, the results confirm that the conv7 structure consistently outperforms the conv3 configuration before applying $TTA^{2}$. On the CS with 10-second audio, the accuracy improved from 47.57% (conv3) to 65.37% (conv7) for single-view augmentation, and from 55.98% to 68.87% when using four augmented views. For SC1 with 1-second inputs, the conv7 kernel achieved 95.93% compared to 96.18% from the smaller conv3 kernel. Although the performance difference for SC1 was marginal, the results show that conv7 maintains or improves accuracy while providing better temporal representation for longer audio sequences.
Overall, these results suggest that medium-sized kernels (such as conv7) provide a good balance between short and long audio lengths, yielding stable performance for the 1D CNN bottleneck in AMAuT.
CLS+TAL Tokens
In AMAuT, the CLS (class) and TAL (tail) tokens are positioned at opposite ends of the token sequence, rather than adjacent to each other as in conventional Transformer designs. This placement allows the model to capture complementary contextual information from both the beginning and end of the sequence, enhancing the overall representational diversity during attention processing.
Table 5: Comparing different positional embedding strategies on CochlScene (CS) and VocalSound (VS) before applying $TTA^{2}$. The tested configurations include CLS+TAL (proposed), CLS+Distil [35], and CLS only [11]. ’Length’ denotes audio length in seconds.
Set Length Method Accuracy (%)
CS 10 s CLS+TAL 68.87
CS 10 s CLS+Distil 69.49
CS 10 s CLS 67.13
VS 12 s CLS+TAL 91.06
VS 12 s CLS+Distil 90.61
VS 12 s CLS 90.56
The comparative results summarized in Table 5 demonstrate that the CLS+TAL positional embedding mechanism introduced in this study serves as an alternative to the commonly used CLS+Distil embedding configuration proposed in DeiT [35].
Contributions of $TTA^{2}$
The four refinement strategies, such as Augmentation (Aug), Multi-Training (Mlt), Hybrid (Hyb), and Test-Time Domain Adaptation (TTDA), each aim to improve AMAuT’s inference robustness and overall testing accuracy. These techniques operate without additional labeled data and are designed to enhance the model’s generalization to unseen domains.
Table 6: Test accuracy improvement (%) from different refinement strategies, including Augmentation (Aug), Multi-training (Mlt), Hybrid (Hyb), and Test-time Domain Adaptation (TTDA), applied to AMAuT. For AudioMNIST (AM), results are reported over five folds (AM0–AM4).
Set Orig Aug Mlt Hyb TTDA
[…]
Experimental results (Table 6) show that all four methods improved test performance on datasets such as SC1, SC2, VS, and CS. For instance, in SC1, accuracy increased from 95.93% to 96.31% with the Hybrid method, while TTDA achieved 96.27%, showing that both ensemble-style and adaptive refinement approaches are effective. Similar improvements were observed for SC2, with accuracy rising from 95.76% to 96.21% using Mlt and 96.13% with TTDA.
For the AudioMNIST (AM) dataset, the effects of the four refinement methods were not entirely consistent. While TTDA and Mlt achieved slightly higher cross-validation accuracies (99.82% and 99.75%, respectively) compared to the baseline (99.71%), the Aug and Hyb methods produced minor decreases (99.66% and 99.45%, respectively).
Despite these variations, the results demonstrate that applying $TTA^{2}$ can still yield incremental benefits depending on dataset characteristics and data distributions.
Table 7: Adaptation time comparison for different $TTA^{2}$ methods in AMAuT using a single RTX 4090 GPU. Time is shown in seconds (s) or minutes (m) per method. “TTDA-Epoch” denotes time costing and the number of epochs used for performing TTDA.
Set Aug Mlt Hyb TTDA-Epoch
SC1 8 s 8 s 24 s 16 m-50
SC2 11 s 11 s 31 s 40 m-80
AM 8s 7 s 20 s 14 m-50
VS 9s 8 s 20 s 16 m-50
CS 64s 23 s 65 s 59 m-50
In terms of computational efficiency, the Mlt method required the least inference time, between 7 and 23 seconds, as reported in Table 7. By comparison, TTDA was more time-consuming because it involves iterative adaptation across multiple epochs (e.g., 50–80 epochs, lasting 14–59 minutes, depending on the dataset).
In summary, the results confirm that all four $TTA^{2}$ strategies can positively impact AMAuT’s generalization at inference. Their effectiveness varies across datasets, suggesting that dataset characteristics, computational constraints, and desired inference efficiency should guide the choice of refinement method. These findings also highlight a promising direction for future research: developing adaptive strategies to automatically select $TTA^{2}$ methods for optimal performance across different domain settings.
4.3 Comparative Performance
Table 8: Comparison of prediction accuracy (%) between the proposed AMAuT and state-of-the-art pre-trained models on five benchmark datasets. Performance values for baseline models are obtained from their respective original publications.
Algorithm Set Accuracy(%)
AMAuT SC1 96.31
SSAST [16] SC1 96.20
AMAuT SC2 96.21
EAT [6] SC2 98.30
AST [14] SC2 98.11
PDC [7] SC2 97.8
AMAuT AM 99.82
HuBERT [19, 27] AM 99.95
XLS-R [2, 27] AM 99.86
Wav2Vec2 [3, 27] AM 99.42
AMAuT VS 92.26
Qwen-Audio [8] VS 92.89
EfficientNet [17, 34] VS 90.5
AMAuT CS 72.58
Audio Flamingo [25] CS 83.0
Qwen-Audio [8] CS 79.5
Table 8 compares AMAuT with several state-of-the-art pre-trained models across the five datasets. AMAuT achieved 96.31% on SC1 and 99.82% on AM, closely matching or exceeding prior results, while maintaining minimal computational overhead.
On VS, AMAuT’s accuracy (92.26%) was only 0.63% below Qwen-Audio [9], and it remained competitive even on the challenging CS dataset (72.58%), where pre-trained models still hold an advantage.
The results suggest that, with an adaptable design, AMAuT trained entirely from scratch can approach or match pre-trained performance in many scenarios.
4.4 Computational Efficiency
By removing the dependence on pre-training, AMAuT achieves remarkable computational efficiency. The training of AMAuT can be completed entirely on a single NVIDIA RTX 4090 GPU, while maintaining performance comparable to or better than several leading pre-trained models such as SSAST [16], EAT [6], HuBERT [19, 27], and Qwen-Audio [8] (see Table 9).
Table 9: Comparison of pre-training and fine-tuning hours for AMAuT and state-of-the-art pre-trained models across multiple datasets. Pre-training and fine-tuning durations are reported for each algorithm, where available, from their respective original publications.
Algorithm Set Pre-train Fine-Tuning
AMAuT SC1 N/A 4 h 10 m
SSAST [16] SC1 10 days [16] unreported
AMAuT SC2 N/A 5 h 38 m
EAT [6] SC2 58h [6] unreported
AST [14] SC2 3 days [35] unreported
PDC [7] SC2 unreported unreported
AMAuT AM N/A 26 m
HuBERT [19, 27] AM 137.75 h [19] unreported
XLS-R [2, 27] AM unreported unreported
Wav2Vec2 [3, 27] AM 7.5 days [3] unreported
AMAuT VS N/A 1 h 28 m
Qwen-Audio [8] VS unreported unreported
EfficientNet [17, 34] VS unknown unreported
AMAuT CS N/A 4 h 9 m
Audio Flamingo [25] CS unreported unreported
Qwen-Audio [8] CS unreported unreported
As shown in Table 9, SSAST [19] requires four NVIDIA GTX Titan X GPUs, 1.95 million AudioSet [13] samples, and 281 thousand LibriSpeech [30] samples for pre-training, a process that takes approximately 10 days. EAT [6] performs pre-training on four RTX 3090 GPUs using the AudioSet [13], requiring about 58 hours. AST [14] does not perform self-pre-training but instead reuses weights from DeiT [35], which itself was trained on ImageNet [32] using four 16 GB V100 GPUs for three days. HuBERT [19, 27] needs 32 GPUs and a total of 60,960 hours of audio (960 hours from LibriSpeech [30] plus 60,000 hours from Libri-Light [22]) for pre-training, which takes approximately 137.75 hours. Similarly, XLS-R [2, 27] requires 128 GPUs and several large-scale multilingual datasets, including VoxPopuli [40], Multilingual LibriSpeech [31], CommonVoice [1], VoxLingua107 [37], and BABEL [12]. Wav2Vec 2.0 [3, 27] also needs 128 V100 GPUs, running for 2.3 days on LibriSpeech [30] and 5.2 days on Libri-Light [22].
Although the pre-training durations and hardware requirements of Qwen-Audio [8] are not explicitly reported, its large-scale use of 30 audio datasets (covering speech, sound, and music) suggests that its pre-training would also be highly resource-intensive. Likewise, Audio Flamingo [25] indicates the use of eight NVIDIA A100 GPUs for pre-training but does not disclose the total compute time or energy cost.
Table 10: Comparison of GPU hours for AMAuT and state-of-the-art pre-trained models across multiple datasets. GPU hours = number of GPUs $\times$ (pre-training + training time). Pre-training and fine-tuning durations are reported for each algorithm where available. Data are collected from original publications unless otherwise noted.
Set AMAuT Baseline Baseline-hour Ratio (%)
SC1 4 h 10 m SSAST $4\times 10$ days 0.43
SC2 5 h 38 m EAT $4\times 58$ h 2.43
SC2 5 h 38 m AST $4\times 3$ days 1.96
AM 26 m HuBERT $32\times 137.75$ h 0.01
AM 26 m Wav2Vec2 $128\times 7.5$ days <0.01<0.01
For GPU hour comparison, AMAuT GPU-hours = per-dataset training time on a single RTX 4090. Baseline GPU-hours are computed as (number of GPUs × reported pretraining hours), where both elements are reported in the baseline paper; if one element is missing, ignore it. Therefore, Table 10 compares AMAuT fine-tuning/training cost (single GPU) to baseline pretraining GPU-hours as reported in the literature; it does not compare training/pre-training times under identical hardware or consider baseline fine-tuning costs. Across all datasets, AMAuT required no more than 3% of the GPU hours of these pre-trained counterparts (see Table 10).
In summary, AMAuT achieves accuracy comparable to or exceeding that of state-of-the-art pre-trained models, while operating at a fraction of their GPU hours. This efficiency highlights the practicality of training AMAuT from scratch on a single high-end GPU, making it particularly attractive for researchers and practitioners with limited computational resources.
5 Discussion
This section discusses the major insights drawn from the study.
Efficiency
The most striking outcome is that AMAuT achieves comparable accuracy to the state-of-the-art across all five benchmarks while requiring less than 3% of the GPU hours of pre-trained models. This result indicates that the architectural adaptability of AMAuT can partially substitute for large-scale pre-training, especially in domain-specific applications.
Flexibility
AMAuT works across sample rates 16–48 kHz and durations 1–12 s without retraining. For 1D CNN, medium-sized kernels (such as conv7) provide a good balance between short and long audio lengths, yielding stable performance for the 1D CNN bottleneck in AMAuT.
Robustness
Augmentation-driven multiview learning substantially improves robustness over traditional sequential augmentation. Because each epoch generates distinct perturbations, the model continually encounters new audio variants, enhancing its generalization to unseen environments. This finding suggests that online multiview augmentation can be a practical alternative to massive data augmentation pipelines.
Further, $TTA^{2}$ improves mean accuracy by 0.4–2% depending on datasets.
6 Limitations and Future Work
Despite its promising results, several limitations remain:
  1. 1.
    Dataset Size Sensitivity
    For stable convergence, datasets with fewer than 15,000 samples tend to underperform, suggesting a lower bound for training purely from scratch.
  2. 2.
    Hyperparameter Dependence
    Performance is sensitive to the Mel-spectrogram [36, 20] configuration and TTDA loss parameters (Eq. 5). The absence of universal tuning guidelines indicates a need for automated hyperparameter search in future work.
  3. 3.
    Inference Latency in TTDA
    While efficient, TTDA still adds 15–60 minutes per dataset during adaptation, limiting real-time deployment (see Table 7).
Future Directions:
  • Investigate long-sequence modeling (>>30 seconds) using hierarchical temporal segmentation to maintain efficiency.
  • Developing hyperparameter search methods for Mel-spectrogram and TTDA loss.
  • Enabling effective training on small-scale datasets with fewer than 15,000 samples.
7 Conclusion
This paper presents AMAuT, an energy-efficient transformer framework for audio classification, designed as a practical alternative to large foundational models. By eliminating the dependency on pre-training, AMAuT overcomes the limitations of fixed sample rates and restricted audio lengths inherent in existing models. To achieve this, this study introduced a novel architecture and training methodology incorporating augmentation-driven multiview learning, horizontal tokenization, a 1D CNN bottleneck, CLS+TAL tokens, and test-time augmentation/adaptation mechanisms ($TTA^{2}$).
Consequently, AMAuT can be trained from scratch on diverse datasets. Evaluations on five benchmark datasets, such as AM, SC1, SC2, VS, and CS, demonstrate that AMAuT achieves performance comparable to state-of-the-art pre-trained models like SSAST [16], EAT [6], HuBERT [19], Qwen-Audio [8], and Audio Flamingo [25], with test accuracies of 99.82%, 96.31%, 96.21%, 92.26%, and 72.58%, respectively. Notably, this performance is achieved using under 3% of the GPU time required by the compared models.
AMAuT demonstrates that pretrain-free Transformers can achieve state-of-the-art performance with drastically reduced computational cost. This efficiency opens new possibilities for practical deployment in embedded, real-time, and low-resource environments.
Acknowledgments
The authors gratefully acknowledge the use of language-editing tools to enhance readability. All interpretations and conclusions remain solely those of the authors.
Data and Code Availability
The AMAuT implementation and experiment scripts are publicly available at https://github.com/Andy-Shao/AMAuT to support full reproducibility of results.
Appendix A Algorithm Details
A.1 Multiview Augmentation
Algorithm 1 Augmentation-driven Multiview Learning
0: ($X,Y$) is the sample features and labels. $lr>0$ is the learning rate.
[…]
1: random initialize $\theta$
2:for $X_{i},Y_{i}\sim\mathcal{U}(X,Y)$ do
3:  $L\leftarrow 0$
4:  for $k\leftarrow 1$ to $K$ do
5:   $\bar{X}_{i}\leftarrow aug_{k}(X_{i})$
6:   $L\leftarrow L+\mathcal{L}(f(\bar{X}_{i};\theta),Y_{i})$
7:  end for
8:  $\theta\leftarrow\theta-lr\nabla_{\theta}L$
9:end for
A.2 Test-time Domain Adaptation (TTDA)
The three components of TTDA are defined as follows:
  • Nuclear-Norm Maximization (NM) [26]:
    Encourages diverse class separation.
[…]
  • Entropy Minimization (EN) [41]:
    Generally, a low entropy value indicates a high degree of certainty. Reducing entropy is reducing uncertainty.
    $\mathcal{L}_{EN}=-\frac{1}{B}\sum^{B}_{j}\sum_{i}^{C}\hat{p}_{i,j}\log{(\hat{p}_{i,j}+\epsilon)}$ (8)
  • A modified Generalized Entropy (GEN) [23]:
    $\mathcal{L}_{GEN}=\frac{1}{B}\sum^{B}_{j}\frac{1-\sum_{i}^{C}\hat{p}^{q}_{i,j}}{q-1}$ (9)
Here, $B$ denotes the batch size, $C$ the number of classes, $q$ is scaling coefficient, $\epsilon=10^{-6}$ avoids division by zero, and $||\cdot||_{F}$ the Frobenius norm.
A.3 Test-time Augmentation (TTAu)
(a) Augmentation-driven refinement (Aug)
This common TTAu approach generates $A$ (even number) weakly augmented variants of each test input and averages their predictions with that of the original sample:
$\begin{split}&\text{aug\_rf}(f,x,A;\theta)\\ &=\frac{1}{A+1}\Big[f(x;\theta)+\sum_{i=1}^{A}f(\text{aug}^{(i)}(x);\theta)\Big]\end{split}$ (10)
where $x$ is a test sample, and $aug^{(i)}$ is the $i$-th augmentation function. In experiments, Aug applied three views, such as (1) the original sample $x$, (2-3) left and right time-shift augmentations ($\pm 17\%$), consistent with those used during prediction, to average outputs. All augmentations were executed on the CPU, so processing time depends on the efficiency of CPU-GPU coordination.
(b) Multi-training-based refinement (Mlt)
This strategy averages predictions from multiple independently trained AMAuT models:
$\text{mlt}(x)=\frac{1}{M}\sum_{i=1}^{M}f(x;\theta^{(i)})$ (11)
where $M$ is an odd integer ($M=2n+1,n\in\mathbb{Z}$) and $\theta^{(i)}$ denotes the parameters of the $i$-th trained model. In experiments, $M=3$. This ensemble-like refinement mitigates model-specific biases.
(c) Hybrid refinement (Hyb)
The hybrid method combines the benefits of both augmentation- and multi-training-based refinements:
$\text{hyb}(x)=\frac{1}{M}\sum_{i=1}^{M}\text{aug\_rf}(f,x,A;\theta^{(i)})$ (12)
In experiments, $M=3$ and $A=2$.
A.4 Training Loss
To prevent overconfidence and improve generalization, this study leverages cross-entropy and label-smoothing regularization (LSR) [33] for training loss:
$\mathcal{L}_{Tr}=-\frac{1}{B}\sum^{B}_{j}\sum^{C}_{i}\big[(1-\epsilon)p_{i,j}+\frac{\epsilon}{C}\big]\log{\hat{p}_{i,j}}$ (13)
where $B$ is the batch size, $C$ the number of classes, $\epsilon=0.1$ the smoothing parameter, $p_{i,j}$ the ground-truth probability, and $\hat{p}_{i,j}$ the model’s predicted probability for class $i$ for sample $j$. This formulation improves model stability and reduces overfitting during training.
Appendix B Hyperparameter Settings
B.1 TTDA Hypermeters
Table 11: Test-time domain adaptation hyperparameter settings
Set $\alpha$ $\beta$ $\gamma$ q
AM 1.0 0.5 0.5 1.1
SC1 0.2 1.0 0.0 0.8
SC2 0.0 1.0 0.2 1.1
VS 1.0 0.5 0.5 1.1
CS 1.0 0.5 0.5 1.1
B.2 Learning Rate Scheduler Hyperparameters
Table 12: Learning rate scheduler hyperparameter settings
Set lr $\lambda$ $\eta$
AM $10^{-3}$ 10 40
SC1 $10^{-3}$ 10 50
SC2 $10^{-3}$ 10 50
VS $10^{-3}$ 10 80
CS $10^{-3}$ 10 40

### v2

@@
 Audio classification is a machine learning task that aims to identify semantic categories from audio signals.
-Pre-trained models such as SSAST [16], EAT [6], HuBERT [19], Qwen-Audio [8], and Audio Flamingo [25], have achieved state-of-the-art performance on benchmark datasets including SpeechCommands V1 (SC1) [42], SpeechCommands V2 (SC2) [42], AudioMNIST (AM) [4], VocalSound (VS) [17, 15], and CochlScene (CS) [21].
+Pre-trained models such as SSAST [gong2022ssast], EAT [chen2024eat], HuBERT [hsu2021hubert], Qwen-Audio [chu2023qwen], and Audio Flamingo [kong2024audio], have achieved state-of-the-art performance on benchmark datasets including SpeechCommands V1 (SC1) & V2 (SC2) [warden2018speech], AudioMNIST (AM) [audiomnist2023], VocalSound (VS) [gong_vocalsound], and CochlScene (CS) [jeong2022cochlscene].
 However, these models rely heavily on pre-training to reach such performance levels, which imposes several practical limitations.
@@
 Although resampling techniques can be applied to meet the required rates, such preprocessing often introduces information loss when downsampling or artificial data when upsampling.
-For instance, [29] reported that the main frequency of bird sounds is below 12 kHz, implying that, based on the information theory [39] (Nyquist theorem), the sample rate should exceed 24 kHz to preserve information.
+For instance, [loo2025temporal] reported that the main frequency of bird sounds is below 12 kHz, implying that, based on the information theory [volkenstein2009entropy] (Nyquist theorem), the sample rate should exceed 24 kHz to preserve information.
 However, 24 kHz is not a standard rate supported by most existing pre-trained models.
@@
 The same limitation applies to audio length.
-SSAST and EAT, for example, support only predefined durations of 1, 5, or 10 seconds [16, 6], while Audio Flamingo allows a maximum of 33.25 seconds [25].
+SSAST and EAT, for example, support only predefined durations of 1, 5, or 10 seconds [gong2022ssast, chen2024eat], while Audio Flamingo allows a maximum of 33.25 seconds [kong2024audio].
 When the input audio length or sample rate does not match the model’s configuration, pre-training must be repeated or the audio must be forcibly adjusted, both of which are computationally expensive.
@@
   3. 3.
-    Dual CLS + TAL tokens: a novel token pairing that captures bidirectional contextual information within the Transformer encoder.
-  4. 4.
     Test-Time adaptation/augmentation ($TTA^{2}$): an inference strategy combining entropy-based domain adaptation and ensemble-style augmentation for improved robustness.
@@
 The transformer architecture, initially developed for language processing, has been effectively adapted for vision and audio tasks.
-Vision Transformer (ViT) [11] and Distilled ViT (DeiT) [35] operate by dividing images into fixed-size patches (e.g., $16\times 16$ or $32\times 32$), a concept later transferred to the audio domain, such as Audio Spectrogram Transformer (AST) [14].
-AST represents audio as Mel-spectrograms [36, 20], turning temporal signals into two-dimensional time–frequency grids before applying patch-based Transformer processing with fixed-size patches ($16\times 16$).
-While AST reuses ImageNet-trained weights [9, 14], SSAST [16] processes pre-training on AudioSet [13] and LibriSpeech [30] rather than reusing ImageNet-trained weights for extending AST.
-Models such as Wav2Vec 2.0 [3] and HuBERT [10] process raw wavforms using 1D CNN-Transformer architectures.
-Wav2Vec 2.0 is composed of seven convolutional layers equipped with LayerNorm and GELU activations, and its large-scale pre-training on LibriSpeech [30] and Libri-Light [22] reportedly required about 7.5 days on 128 GPUs.
-HuBERT follows a similar CNN-Transformer architecture to Wav2Vec 2.0 but extends it with a BERT-style [10] masked prediction mechanism that integrates a waveform encoder.
+Vision Transformer (ViT) [dosovitskiy2020image] and Distilled ViT (DeiT) [touvron2021training] operate by dividing images into fixed-size patches (e.g., $16\times 16$ or $32\times 32$), a concept later transferred to the audio domain, such as Audio Spectrogram Transformer (AST) [gong2021ast].
+AST represents audio as Mel-spectrograms [ustubioglu2023mel, hwang2020mel], turning temporal signals into two-dimensional time–frequency grids before applying patch-based Transformer processing with fixed-size patches ($16\times 16$).
+While AST reuses ImageNet-trained weights [deng2009imagenet, gong2021ast], SSAST [gong2022ssast] processes pre-training on AudioSet [gemmeke2017audio] and LibriSpeech [panayotov2015librispeech] rather than reusing ImageNet-trained weights for extending AST.
+Models such as Wav2Vec 2.0 [baevski2020wav2vec] and HuBERT [devlin2019bert] process raw wavforms using 1D CNN-Transformer architectures.
+Wav2Vec 2.0 is composed of seven convolutional layers equipped with LayerNorm and GELU activations, and its large-scale pre-training on LibriSpeech [panayotov2015librispeech] and Libri-Light [kahn2020libri] reportedly required about 7.5 days on 128 GPUs.
+HuBERT follows a similar CNN-Transformer architecture to Wav2Vec 2.0 but extends it with a BERT-style [devlin2019bert] masked prediction mechanism that integrates a waveform encoder.
 Like Wav2Vec 2.0, AMAuT adopts a 1D CNN-Transformer structure but differs fundamentally: AMAuT trains from scratch without pre-training, uses 1D CNNs on Mel-spectrogram horizontal tokens, and replaces LayerNorm + GELU with BatchNorm + ReLU.
@@
 99.4% on AudioMNIST) between Wav2Vec 2.0 and AMAuT.
-The spectrogram transformers [43] contributed a Tempo-Frequency sequential (TFS) attention for separating mel-spectrogram outputs within frequency direction instead of $16\times 16$ regions in AST [14] that are similar to the horizontal tokenization of AMAuT (see subsection 3.3).
-However, [43] inherits the methodology of AST and ViT, and then [43] employs a different model architecture, as it incorporates 2D CNN layers.
+The spectrogram transformers [zhang2022spectrogram] contributed a Tempo-Frequency sequential (TFS) attention for separating mel-spectrogram outputs within frequency direction instead of $16\times 16$ regions in AST [gong2021ast] that are similar to the horizontal tokenization of AMAuT (see subsection 3.3).
+However, [zhang2022spectrogram] inherits the methodology of AST and ViT, and then [zhang2022spectrogram] employs a different model architecture, as it incorporates 2D CNN layers.
 Overall, both AST and SSAST constrain input to fixed sample rates (typically 16 kHz) and durations (1 s, 5 s, or 10 s) when reusing their pre-trained weights; Wav2Vec 2.0 and HuBERT are computationally expensive and also fixed at 16 kHz.
@@
 2.2 Test-Time Domain Adaptation
-Test-time domain adaptation (TTDA) only uses the test set for performance improvement by adapting (partial) model parameters of a trained model from a source domain to a target domain [28].
+Test-Time domain adaptation (TTDA) only uses the test set for performance improvement by adapting (partial) model parameters of a trained model from a source domain to a target domain [liang2024comprehensive].
 In AMAuT, this is achieved through a composite loss function $\mathcal{L}_{TTDA}$ that combines three complementary objectives: Nuclear-Norm Maximization, Entropy Minimization, and Generalized Entropy (see Eq.7, Eq.8, and Eq.9 in Appendix A).
-These components have been applied separately in previous studies, such as CoNMix [26] for test-time domain adaptation, TENT [41] for entropy-based test-time batch adaptation, and SGEM [23] for Automatic Speech Recognition, but AMAuT unifies them for the first time.
+These components have been applied separately in previous studies, such as CoNMix [kumar2023conmix] for test-time domain adaptation, TENT [wang2021tent] for entropy-based test-time batch adaptation, and SGEM [kim2023sgem] for Automatic Speech Recognition, but AMAuT unifies them for the first time.
 This integration allows efficient test-time domain adaptation and improves robustness under moderate distribution shifts without additional supervision.
@@
 Specifically:
-  • Training (four augmentations): (1) random left or right time shift (0-17%), (2) Gaussian noise (noise-to-audio ratio = 0.015, and regenerate noise each time), and (3-4) random clips from two background noises (“dude miaowing” and “pick noise” from SpeechCommands [42]) with an SNR of 50 dB.
+  • Training (four augmentations): (1) random left or right time shift (0-17%), (2) Gaussian noise (noise-to-audio ratio = 0.015, and regenerate noise each time), and (3-4) random clips from two background noises (“dude miaowing” and “pick noise” from SpeechCommands [warden2018speech]) with an SNR of 50 dB.
   • TTDA (two augmentations): (1) random right time shift (0-17%), and (2) random left time shift (0-17%).
@@
 3 and Figure 3).
-Unlike conventional designs such as DeiT [35], where CLS and distillation tokens are placed adjacently, AMAuT positions CLS and TAL at opposite ends of the token sequence.
-This design assumes that tokens at different spatial positions may encode distinct contextual information during transformer processing.
+Similar to DeiT [touvron2021training], AMAuT positions CLS and TAL at opposite ends of the token sequence.
 Before the attention extraction stage, AMAuT applies a learnable positional embedding and a random dropout layer (15% drop rate) to the concatenated token sequence:
@@
 Attention Extraction
-This component follows the Transformer [38] architecture and comprises multiple transformer blocks. Two important configurations are adopted:
+This component follows the Transformer [vaswani2017attention] architecture and comprises multiple transformer blocks. Two important configurations are adopted:
   1. 1.
-    The dropout rate for both multi-head attention and MLP is set to 0%, consistent with models such as SSAST [16], AST [14], and CoNMix [26].
+    The dropout rate for both multi-head attention and MLP is set to 0%, consistent with models such as SSAST [gong2022ssast], AST [gong2021ast], and CoNMix [kumar2023conmix].
   2. 2.
@@
 This subsection investigates an uncommon application of Test-time Domain Adaptation (TTDA): improving accuracy when training and test distributions are similar.
-Typically, TTDA is employed to address distribution shifts, but we explore its utility even when such a shift is minimal [28].
-TTDA enhances performance by adapting a trained model to the test data distribution without requiring labels [28].
+Typically, TTDA is employed to address distribution shifts, but we explore its utility even when such a shift is minimal [liang2024comprehensive].
+TTDA enhances performance by adapting a trained model to the test data distribution without requiring labels [liang2024comprehensive].
 During TTDA, all test features are visible, but labels remain hidden.
@@
 $\mathcal{L}_{TTDA}=\alpha\mathcal{L}_{NM}+\beta\mathcal{L}_{EN}+\gamma\mathcal{L}_{GEN}$ (5)
-where $\alpha$, $\beta$, and $\gamma$ are scaling coefficients (see Table11 in Appedix B).
+where $\alpha$, $\beta$, and $\gamma$ are scaling coefficients (see Table12 in Appedix C).
 By integrating these complementary objectives, AMAuT improves inference robustness in scenarios where training and testing data share similar but not identical distributions.
 Test-Time Augmentation (TTAu)
-To further enhance inference robustness, AMAuT adopts test-time augmentation (TTAu) [24].
+To further enhance inference robustness, AMAuT adopts test-time augmentation (TTAu) [kimura2021understanding].
 Unlike TTDA, TTAu does not update model parameters; instead, it refines predictions by averaging outputs from multiple augmented versions of the same input.
@@
 where $e$ denotes the epoch number, $10\leq\lambda\leq 40$ controls the rate of change, $200\geq\eta\geq 40$ defines the cardinality.
-See more details of hyperparameter settings from Table 12 in Appendix B.
+See more details of hyperparameter settings from Table 12 in Appendix C.
 During both training and TTDA, AMAuT uses the stochastic gradient descent (SGD) optimizer with weight decay $10^{-3}$, momentum 0.9, and Nesterov acceleration enabled.
 4 Experiments and Results
-AMAuT integrates several architectural and training strategies, including augmentation-driven multiview learning, conv1 + conv7 + conv1 CNN bottlenecks, flexible handling of arbitrary sample rates and audio lengths, CLS+TAL learnable positional embeddings, and $TTA^{2}$.
+AMAuT integrates several architectural and training strategies, including augmentation-driven multiview learning, conv1 + conv7 + conv1 CNN bottlenecks, flexible handling of arbitrary sample rates and audio lengths, learnable positional embeddings, and $TTA^{2}$.
 Collectively, these components enable AMAuT to achieve strong performance when trained entirely from scratch, without relying on large-scale pre-training.
@@
 CS 76115 10 s 44.1 kHz
-To evaluate the proposed AMAuT framework, this study employed five publicly available benchmark datasets: AudioMNIST (AM) [4], SpeechCommands V1 (SC1) [42], SpeechCommands V2 (SC2) [42], VocalSound (VS) [17, 15], and CochlScene (CS) [21].
+To evaluate the proposed AMAuT framework, this study employed five publicly available benchmark datasets: AudioMNIST (AM) [audiomnist2023], SpeechCommands V1 (SC1) & V2 (SC2) [warden2018speech], VocalSound (VS) [gong_vocalsound], and CochlScene (CS) [jeong2022cochlscene].
 As shown in Table 2, these datasets cover a diverse range of sampling rates and audio durations, enabling a comprehensive analysis of AMAuT’s flexibility and robustness.
@@
 For model training, AMAuT used the training and validation splits for AM, and the training sets for SC1, SC2, VS, and CS.
+4.2 Comparative Performance
+Table 3 compares AMAuT with several state-of-the-art pre-trained models across the five datasets.
+Table 3: Comparison of prediction accuracy (%) between the proposed AMAuT and state-of-the-art pre-trained models on five benchmark datasets.
+Algorithm Set Accuracy(%)
+AMAuT SC1 96.31
+SSAST [gong2022ssast] SC1 96.20
+AMAuT SC2 96.21
+EAT [chen2024eat] SC2 98.30
+AST [gong2021ast] SC2 98.11
+PDC [chrysos2022augmenting] SC2 97.8
+AMAuT AM 99.82
+HuBERT [hsu2021hubert, la2024benchmarking] AM 99.95
+XLS-R [babu2021xls, la2024benchmarking] AM 99.86
+Wav2Vec2 [baevski2020wav2vec, la2024benchmarking] AM 99.42
+AMAuT VS 92.26
+Qwen-Audio [chu2023qwen] VS 92.89
+EfficientNet [gong_vocalsound, tan2019efficientnet] VS 90.5
+AMAuT CS 72.58
+Audio Flamingo [kong2024audio] CS 83.0
+Qwen-Audio [chu2023qwen] CS 79.5
+  • 1
+    Performance values for baseline models are obtained from their respective original publications.
+AMAuT achieved 96.31% on SC1 and 99.82% on AM, closely matching or exceeding prior results, while maintaining minimal computational overhead.
+On VS, AMAuT’s accuracy (92.26%) was only 0.63% below Qwen-Audio [9], and it remained competitive even on the challenging CS dataset (72.58%), where pre-trained models still hold an advantage.
+The results suggest that, with an adaptable design, AMAuT trained entirely from scratch can approach or match pre-trained performance in many scenarios.
 4.3 Ablation Studies
@@
 Consequently, the model is continually exposed to diverse variants of the same source sample, improving generalization and robustness.
-Table 3: Comparison of test accuracies between sequential augmentation and augmentation-driven multiview learning on CochlScene (CS) and SpeechCommands V1 (SC1) before applying $TTA^{2}$.
-“1-Aug” and “4-Aug” indicate the use of one or four augmented views, respectively.
-The arrows ($\uparrow$, $\downarrow$) denote performance improvement or degradation relative to the baseline.
+Table 4: Comparison of test accuracies between sequential augmentation and augmentation-driven multiview learning on CS and SC1 before applying $TTA^{2}$.
 Set Aug-method 1-Aug(%) 4-Aug(%)
@@
 […]
+  • 1
+    “1-Aug” and “4-Aug” indicate the use of one or four augmented views, respectively.
+  • 2
+    The arrows ($\uparrow$, $\downarrow$) denote performance improvement or degradation relative to the baseline.
 Compared with the traditional dataset-extension method (offline), where augmented samples are pre-generated and stored for later training, the augmentation-driven multiview learning approach offers two significant advantages:
@@
 For online augmentation methods, as shown in Table 4, the proposed multiview approach achieved higher test accuracies on both datasets, for example, an increase from 65.37% to 68.87% on CS and from 95.26% to 95.93% on SC1, demonstrating that real-time, multiview training substantially improves performance without additional pre-training or data storage requirements.
+Consistency Analysis:
+To further verify that augmentation-driven multiview learning improves convergence stability and generalization, a consistency analysis was conducted using the agreement rate metric (see Eq.
+15 in Appendix B).
+Table 5: Comparison of agreement rates between one and four augmentation-driven multiview configurations for AMAuT across five benchmark datasets before applying $TTA^{2}$.
+Set Length 1-Aug 4-Aug
+[…]
+[…]
+[…]
+[…]
+AM 1 s 0.9995 N/A
+  • 1
+    “1-Aug” and “4-Aug” indicate the use of one or four augmented views, respectively.
+  • 2
+    Agreement rate ranges from 0 to 1, where 1 denotes perfect consistency.
+  • 3
+    For training AM, multiple views are unnecessary, as a single view already achieves above 99.25% accuracy (see Table 7).
+As shown in Table 5, the four-view configuration consistently yielded higher agreement rates across CS, VS, SC1, and SC2.
+This improvement indicates that augmentation-driven multiview learning enhances model stability and convergence reliability.
+The results thus provide empirical evidence that multiview augmentation effectively strengthens the robustness and reproducibility of the AMAuT framework.
 CNN Kernel Design
 In many existing CNN-based audio classification architectures, relatively small convolutional kernels are commonly used.
-For example, HuBERT [19] and Wav2Vec 2.0 [3] employ bottleneck structures such as conv2 + conv3, while other hybrid CNN-Transformer architectures like ENACT-Heart [18] and CNN-TE [5] use kernel sizes of 3, 2, and 1.
+For example, HuBERT [hsu2021hubert] and Wav2Vec 2.0 [baevski2020wav2vec] employ bottleneck structures such as conv2 + conv3, while other hybrid CNN-Transformer architectures like ENACT-Heart [han2025enact] and CNN-TE [chen2024hybrid] use kernel sizes of 3, 2, and 1.
 These designs typically include only one or two large convolutional layers, such as conv14 or conv16, at the head or tail of the CNN stack, rather than throughout its body.
@@
 The conv7 kernel allows AMAuT to capture longer temporal dependencies without significantly increasing computational complexity.
-Table 4: Performance comparison between conv3 and conv7 kernel structures on CochlScene (CS) and SpeechCommands V1 (SC1) before applying $(TTA)^{2}$, evaluated on a single RTX 4090 GPU.
-“Length” indicates audio length in seconds, and “1-Aug” / “4-Aug” denote the number of augmented views used during testing.
+Table 6: Performance comparison between conv3 and conv7 kernel structures on CS and SC1 before applying $(TTA)^{2}$.
 Set Length Structure 1-Aug(%) 4-Aug(%)
@@
 […]
+  • 1
+    “Length” indicates audio length in seconds.
+  • 2
+    “1-Aug” / “4-Aug” denotes the number of augmented views used during testing.
+  • 3
+    Evaluated on a single RTX 4090 GPU
 As presented in Table 6, the results confirm that the conv7 structure consistently outperforms the conv3 configuration before applying $TTA^{2}$.
@@
 Overall, these results suggest that medium-sized kernels (such as conv7) provide a good balance between short and long audio lengths, yielding stable performance for the 1D CNN bottleneck in AMAuT.
-CLS+TAL Tokens
-In AMAuT, the CLS (class) and TAL (tail) tokens are positioned at opposite ends of the token sequence, rather than adjacent to each other as in conventional Transformer designs.
-This placement allows the model to capture complementary contextual information from both the beginning and end of the sequence, enhancing the overall representational diversity during attention processing.
-Table 5: Comparing different positional embedding strategies on CochlScene (CS) and VocalSound (VS) before applying $TTA^{2}$.
-The tested configurations include CLS+TAL (proposed), CLS+Distil [35], and CLS only [11].
-’Length’ denotes audio length in seconds.
-Set Length Method Accuracy (%)
-CS 10 s CLS+TAL 68.87
-CS 10 s CLS+Distil 69.49
-CS 10 s CLS 67.13
-VS 12 s CLS+TAL 91.06
-VS 12 s CLS+Distil 90.61
-VS 12 s CLS 90.56
-The comparative results summarized in Table 5 demonstrate that the CLS+TAL positional embedding mechanism introduced in this study serves as an alternative to the commonly used CLS+Distil embedding configuration proposed in DeiT [35].
 Contributions of $TTA^{2}$
@@
 These techniques operate without additional labeled data and are designed to enhance the model’s generalization to unseen domains.
-Table 6: Test accuracy improvement (%) from different refinement strategies, including Augmentation (Aug), Multi-training (Mlt), Hybrid (Hyb), and Test-time Domain Adaptation (TTDA), applied to AMAuT.
-For AudioMNIST (AM), results are reported over five folds (AM0–AM4).
+Table 7: Test accuracy improvement (%) from different refinement strategies, including Aug, Mlt, Hyb, and TTDA, applied to AMAuT.
 Set Orig Aug Mlt Hyb TTDA
@@
 […]
+  • 1
+    For AudioMNIST (AM), results are reported over five folds (AM0–AM4).
 Experimental results (Table 7) show that all four methods improved test performance on datasets such as SC1, SC2, VS, and CS.
@@
 Table 8: Adaptation time comparison for different $TTA^{2}$ methods in AMAuT using a single RTX 4090 GPU.
-Time is shown in seconds (s) or minutes (m) per method.
-“TTDA-Epoch” denotes time costing and the number of epochs used for performing TTDA.
 Set Aug Mlt Hyb TTDA-Epoch
@@
 CS 64s 23 s 65 s 59 m-50
+  • 1
+    Time is shown in seconds (s) or minutes (m) per method.
+  • 2
+    “TTDA-Epoch” denotes time costing and the number of epochs used for performing TTDA.
 In terms of computational efficiency, the Mlt method required the least inference time, between 7 and 23 seconds, as reported in Table 8.
@@
 These findings also highlight a promising direction for future research: developing adaptive strategies to automatically select $TTA^{2}$ methods for optimal performance across different domain settings.
-4.3 Comparative Performance
-Table 8: Comparison of prediction accuracy (%) between the proposed AMAuT and state-of-the-art pre-trained models on five benchmark datasets.
-Performance values for baseline models are obtained from their respective original publications.
-Algorithm Set Accuracy(%)
-AMAuT SC1 96.31
-SSAST [16] SC1 96.20
-AMAuT SC2 96.21
-EAT [6] SC2 98.30
-AST [14] SC2 98.11
-PDC [7] SC2 97.8
-AMAuT AM 99.82
-HuBERT [19, 27] AM 99.95
-XLS-R [2, 27] AM 99.86
-Wav2Vec2 [3, 27] AM 99.42
-AMAuT VS 92.26
-Qwen-Audio [8] VS 92.89
-EfficientNet [17, 34] VS 90.5
-AMAuT CS 72.58
-Audio Flamingo [25] CS 83.0
-Qwen-Audio [8] CS 79.5
-Table 8 compares AMAuT with several state-of-the-art pre-trained models across the five datasets.
-AMAuT achieved 96.31% on SC1 and 99.82% on AM, closely matching or exceeding prior results, while maintaining minimal computational overhead.
-On VS, AMAuT’s accuracy (92.26%) was only 0.63% below Qwen-Audio [9], and it remained competitive even on the challenging CS dataset (72.58%), where pre-trained models still hold an advantage.
-The results suggest that, with an adaptable design, AMAuT trained entirely from scratch can approach or match pre-trained performance in many scenarios.
+CLS+TAL Tokens
+In AMAuT, the CLS (class) and TAL (tail) tokens are positioned at opposite ends of the token sequence (Similar to DeiT [touvron2021training]).
+Table 9: Comparing different positional embedding strategies on CS and VS before applying $TTA^{2}$.
+Set Length Method Accuracy (%)
+CS 10 s CLS+TAL 68.87
+CS 10 s CLS 67.13
+VS 12 s CLS+TAL 91.06
+VS 12 s CLS 90.56
+  • 1
+    See CLS token in [dosovitskiy2020image, touvron2021training].
+  • 2
+    ’Length’ denotes audio length in seconds.
+The results summarized in Table 9 show that CLS+TAl tokens perform better than a single CLS token.
 4.4 Computational Efficiency
 By removing the dependence on pre-training, AMAuT achieves remarkable computational efficiency.
-The training of AMAuT can be completed entirely on a single NVIDIA RTX 4090 GPU, while maintaining performance comparable to or better than several leading pre-trained models such as SSAST [16], EAT [6], HuBERT [19, 27], and Qwen-Audio [8] (see Table 9).
+The training of AMAuT can be completed entirely on a single NVIDIA RTX 4090 GPU, while maintaining performance comparable to or better than several leading pre-trained models such as SSAST [gong2022ssast], EAT [chen2024eat], HuBERT [hsu2021hubert, la2024benchmarking], and Qwen-Audio [chu2023qwen] (see Table 10).
 Table 10: Comparison of pre-training and fine-tuning hours for AMAuT and state-of-the-art pre-trained models across multiple datasets.
-Pre-training and fine-tuning durations are reported for each algorithm, where available, from their respective original publications.
 Algorithm Set Pre-train Fine-Tuning
 AMAuT SC1 N/A 4 h 10 m
-SSAST [16] SC1 10 days [16] unreported
+SSAST [gong2022ssast] SC1 10 days [gong2022ssast] unreported
 AMAuT SC2 N/A 5 h 38 m
-EAT [6] SC2 58h [6] unreported
-AST [14] SC2 3 days [35] unreported
-PDC [7] SC2 unreported unreported
+EAT [chen2024eat] SC2 58h [chen2024eat] unreported
+AST [gong2021ast] SC2 3 days [touvron2021training] unreported
+PDC [chrysos2022augmenting] SC2 unreported unreported
 AMAuT AM N/A 26 m
-HuBERT [19, 27] AM 137.75 h [19] unreported
-XLS-R [2, 27] AM unreported unreported
-Wav2Vec2 [3, 27] AM 7.5 days [3] unreported
+HuBERT [hsu2021hubert, la2024benchmarking] AM 137.75 h [hsu2021hubert] unreported
+XLS-R [babu2021xls, la2024benchmarking] AM unreported unreported
+Wav2Vec2 [baevski2020wav2vec, la2024benchmarking] AM 7.5 days [baevski2020wav2vec] unreported
 AMAuT VS N/A 1 h 28 m
-Qwen-Audio [8] VS unreported unreported
-EfficientNet [17, 34] VS unknown unreported
+Qwen-Audio [chu2023qwen] VS unreported unreported
+EfficientNet [gong_vocalsound, tan2019efficientnet] VS unknown unreported
 AMAuT CS N/A 4 h 9 m
-Audio Flamingo [25] CS unreported unreported
-Qwen-Audio [8] CS unreported unreported
-As shown in Table 9, SSAST [19] requires four NVIDIA GTX Titan X GPUs, 1.95 million AudioSet [13] samples, and 281 thousand LibriSpeech [30] samples for pre-training, a process that takes approximately 10 days.
-EAT [6] performs pre-training on four RTX 3090 GPUs using the AudioSet [13], requiring about 58 hours.
-AST [14] does not perform self-pre-training but instead reuses weights from DeiT [35], which itself was trained on ImageNet [32] using four 16 GB V100 GPUs for three days.
-HuBERT [19, 27] needs 32 GPUs and a total of 60,960 hours of audio (960 hours from LibriSpeech [30] plus 60,000 hours from Libri-Light [22]) for pre-training, which takes approximately 137.75 hours.
-Similarly, XLS-R [2, 27] requires 128 GPUs and several large-scale multilingual datasets, including VoxPopuli [40], Multilingual LibriSpeech [31], CommonVoice [1], VoxLingua107 [37], and BABEL [12].
-Wav2Vec 2.0 [3, 27] also needs 128 V100 GPUs, running for 2.3 days on LibriSpeech [30] and 5.2 days on Libri-Light [22].
-Although the pre-training durations and hardware requirements of Qwen-Audio [8] are not explicitly reported, its large-scale use of 30 audio datasets (covering speech, sound, and music) suggests that its pre-training would also be highly resource-intensive.
-Likewise, Audio Flamingo [25] indicates the use of eight NVIDIA A100 GPUs for pre-training but does not disclose the total compute time or energy cost.
+Audio Flamingo [kong2024audio] CS unreported unreported
+Qwen-Audio [chu2023qwen] CS unreported unreported
+  • 1
+    Pre-training and fine-tuning durations are reported for each algorithm, where available, from their respective original publications.
+As shown in Table 10, SSAST [gong2022ssast] requires four NVIDIA GTX Titan X GPUs, 1.95 million AudioSet [gemmeke2017audio] samples, and 281 thousand LibriSpeech [panayotov2015librispeech] samples for pre-training, a process that takes approximately 10 days.
+EAT [chen2024eat] performs pre-training on four RTX 3090 GPUs using the AudioSet [gemmeke2017audio], requiring about 58 hours.
+AST [gong2021ast] does not perform self-pre-training but instead reuses weights from DeiT [touvron2021training], which itself was trained on ImageNet [ILSVRC15] using four 16 GB V100 GPUs for three days.
+HuBERT [hsu2021hubert, la2024benchmarking] needs 32 GPUs and a total of 60,960 hours of audio (960 hours from LibriSpeech [panayotov2015librispeech] plus 60,000 hours from Libri-Light [kahn2020libri]) for pre-training, which takes approximately 137.75 hours.
+Similarly, XLS-R [babu2021xls, la2024benchmarking] requires 128 GPUs and several large-scale multilingual datasets, including VoxPopuli [wang2021voxpopuli], Multilingual LibriSpeech [pratap2020mls], CommonVoice [ardila2019common], VoxLingua107 [valk2021voxlingua107], and BABEL [gales2014speech].
+Wav2Vec 2.0 [baevski2020wav2vec, la2024benchmarking] also needs 128 V100 GPUs, running for 2.3 days on LibriSpeech [panayotov2015librispeech] and 5.2 days on Libri-Light [kahn2020libri].
+Although the pre-training durations and hardware requirements of Qwen-Audio [chu2023qwen] are not explicitly reported, its large-scale use of 30 audio datasets (covering speech, sound, and music) suggests that its pre-training would also be highly resource-intensive.
+Likewise, Audio Flamingo [kong2024audio] indicates the use of eight NVIDIA A100 GPUs for pre-training but does not disclose the total compute time or energy cost.
 Table 11: Comparison of GPU hours for AMAuT and state-of-the-art pre-trained models across multiple datasets.
-GPU hours = number of GPUs $\times$ (pre-training + training time).
-Pre-training and fine-tuning durations are reported for each algorithm where available.
-Data are collected from original publications unless otherwise noted.
 Set AMAuT Baseline Baseline-hour Ratio (%)
@@
 AM 26 m Wav2Vec2 $128\times 7.5$ days <0.01<0.01
+  • 1
+    GPU hours = number of GPUs $\times$ (pre-training + training time).
+  • 2
+    Pre-training and fine-tuning durations are reported for each algorithm where available.
+  • 3
+    Data are collected from original publications unless otherwise noted.
 For GPU hour comparison, AMAuT GPU-hours = per-dataset training time on a single RTX 4090.
@@
     Hyperparameter Dependence
-    Performance is sensitive to the Mel-spectrogram [36, 20] configuration and TTDA loss parameters (Eq.
+    Performance is sensitive to the Mel-spectrogram [ustubioglu2023mel, hwang2020mel] configuration and TTDA loss parameters (Eq.
 5).
@@
 Consequently, AMAuT can be trained from scratch on diverse datasets.
-Evaluations on five benchmark datasets, such as AM, SC1, SC2, VS, and CS, demonstrate that AMAuT achieves performance comparable to state-of-the-art pre-trained models like SSAST [16], EAT [6], HuBERT [19], Qwen-Audio [8], and Audio Flamingo [25], with test accuracies of 99.82%, 96.31%, 96.21%, 92.26%, and 72.58%, respectively.
+Evaluations on five benchmark datasets, such as AM, SC1, SC2, VS, and CS, demonstrate that AMAuT achieves performance comparable to state-of-the-art pre-trained models like SSAST [gong2022ssast], EAT [chen2024eat], HuBERT [hsu2021hubert], Qwen-Audio [chu2023qwen], and Audio Flamingo [kong2024audio], with test accuracies of 99.82%, 96.31%, 96.21%, 92.26%, and 72.58%, respectively.
 Notably, this performance is achieved using under 3% of the GPU time required by the compared models.
@@
 The three components of TTDA are defined as follows:
-  • Nuclear-Norm Maximization (NM) [26]:
+  • Nuclear-Norm Maximization (NM) [kumar2023conmix]:
     Encourages diverse class separation.
 […]
-  • Entropy Minimization (EN) [41]:
+  • Entropy Minimization (EN) [wang2021tent]:
     Generally, a low entropy value indicates a high degree of certainty. Reducing entropy is reducing uncertainty.
     $\mathcal{L}_{EN}=-\frac{1}{B}\sum^{B}_{j}\sum_{i}^{C}\hat{p}_{i,j}\log{(\hat{p}_{i,j}+\epsilon)}$ (8)
-  • A modified Generalized Entropy (GEN) [23]:
+  • A modified Generalized Entropy (GEN) [kim2023sgem]:
     $\mathcal{L}_{GEN}=\frac{1}{B}\sum^{B}_{j}\frac{1-\sum_{i}^{C}\hat{p}^{q}_{i,j}}{q-1}$ (9)
@@
 A.4 Training Loss
-To prevent overconfidence and improve generalization, this study leverages cross-entropy and label-smoothing regularization (LSR) [33] for training loss:
+To prevent overconfidence and improve generalization, this study leverages cross-entropy and label-smoothing regularization (LSR) [szegedy2016rethinking] for training loss:
 $\mathcal{L}_{Tr}=-\frac{1}{B}\sum^{B}_{j}\sum^{C}_{i}\big[(1-\epsilon)p_{i,j}+\frac{\epsilon}{C}\big]\log{\hat{p}_{i,j}}$ (13)
@@
 This formulation improves model stability and reduces overfitting during training.
-Appendix B Hyperparameter Settings
-B.1 TTDA Hypermeters
-Table 11: Test-time domain adaptation hyperparameter settings
-Set $\alpha$ $\beta$ $\gamma$ q
-AM 1.0 0.5 0.5 1.1
-SC1 0.2 1.0 0.0 0.8
-SC2 0.0 1.0 0.2 1.1
-VS 1.0 0.5 0.5 1.1
-CS 1.0 0.5 0.5 1.1
-B.2 Learning Rate Scheduler Hyperparameters
-Table 12: Learning rate scheduler hyperparameter settings
-Set lr $\lambda$ $\eta$
-AM $10^{-3}$ 10 40
-SC1 $10^{-3}$ 10 50
-SC2 $10^{-3}$ 10 50
-VS $10^{-3}$ 10 80
-CS $10^{-3}$ 10 40
+Appendix B Consistency Analysis
+To evaluate the consistency of independently trained AMAuT models, an agreement rate is computed based on the alignment of their predictions.
+Let the confusion matrix between two prediction sets $p_{1}$ and $p_{2}$ be defined as:
+$cm=\text{confusion\_matrix}(p_{1},p_{2})$ (14)
+The overall agreement rate is then calculated as:
+$\text{agreement rate}=\frac{\sum_{i}cm_{i,i}}{\sum_{i,j}cm_{i,j}}$ (15)
+where $cm_{i,i}$ denotes the number of samples for which both predict the same class label $i$, and $\sum_{i,j}cm_{i,j}$ represents the total number of samples considered.
+A higher agreement rate indicates greater consistency between models trained independently under the same configuration, suggesting stable convergence and robust generalization of the AMAuT framework.
+Appendix C Hyperparameter Settings
+Table 12: Hyperparameter Settings
+Set $\alpha$ $\beta$ $\gamma$ $q$ lrlr $\lambda$ $\eta$
+AM 1.0 0.5 0.5 1.1 $10^{-3}$ 10 40
+SC1 0.2 1.0 0.0 0.8 $10^{-3}$ 10 50
+SC2 0.0 1.0 0.2 1.1 $10^{-3}$ 10 50
+VS 1.0 0.5 0.5 1.1 $10^{-3}$ 10 80
+CS 1.0 0.5 0.5 1.1 $10^{-3}$ 10 40
+  • 1
+    The parameters, $\alpha$, $\beta$, and $\gamma$, correspond to the weighting coefficients in Eq. 5.
+  • 2
+    The coefficient $q$ denotes the scaling factor in Eq. 9.
+  • 3
+    The learning rate lrlr, adjustment factor $\lambda$, and scheduling cardinality $\eta$ follow the update rule defined in Eq. 6.